OpenStudy (anonymous):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=8, b=4, c=?

OpenStudy (anonymous):

do u do 8 * 8 and 4 * 4 then what

OpenStudy (anonymous):

\[a^2+b^2=c^2\] \[8^2+4^2=c^2\] \[64+16=c^2\] etc

OpenStudy (anonymous):

so i add those then i square that ?

OpenStudy (anonymous):

add them then take the square root

OpenStudy (anonymous):

so its 80 and then square root that

OpenStudy (anonymous):

yes \(c=\sqrt{80}\)

OpenStudy (anonymous):

and since \(80=16\times 5\) you can write \(c=4\sqrt{5}\)

OpenStudy (anonymous):

i got it :)